OpenStudy (djbreezy):

which is an equation in point-slope form of the line that passes through the points (-4,-1) and (5,7) A. y+4=9/8(x+1) B. Y+1=8/9(x+4) C. Y+1=8/7(x+4) D. Y-1=8/9(x-4)

pooja195 (pooja195):

Use the point slope formula to plug in the given info \[\huge~\rm~y-y_1=m(x-x_1)\]

OpenStudy (mathmale):

To expand on that: You are given two points. Let one of the points (doesn't matter which one) replace (x1, y1) in the above equation, and let the other point replace (x,y). This will enable you to calculate the slope (m) quickly. Now rewrite \[\huge~\rm~y-y_1=m(x-x_1)\], inserting the value of m you've found, and replacing (x1,y1) with either given point. That'll give you the equation of this line in point-slope form.